# Large Language Model Unlearning for Source Code

### v1

@@ Abstract
Large Language Model for Software Engineering (LLM4SE) have demonstrated significant success, but LLMs’ potential memorization of sensitive or outdated training data introduces critical risks to legal compliance, software security, and code quality. LLM unlearning techniques, which can eliminate the influence of undesired data from LLMs in a post-training way, present a promising solution to address these concerns. While recent efforts in LLM unlearning show effectiveness in natural language data, their applicability to source code remains underexplored. Our empirical study reveals that existing LLM unlearning approaches, when applied to source code, cause severe model utility degradation, rendering models practically unusable for code generation. In this paper, we propose PROD, a novel unlearning approach that enables LLMs to forget undesired code content while effectively preserving their code generation capabilities. PROD suppresses the probability of forget data in LLMs’ output distribution while promoting candidate distributional components, enabling the model to jointly learn to forget specific content and retain its general capabilities. To facilitate this study, we establish a benchmark for code unlearning evaluation, which includes three critical downstream tasks: copyrighted code unlearning, insecure code unlearning, and deprecated API unlearning. Our comprehensive evaluation demonstrates that PROD achieves superior balance between forget quality and model utility compared to existing unlearning approaches across three downstream tasks, while consistently exhibiting improvements when applied to LLMs of varying series. PROD also exhibits superior robustness against adversarial attacks without generating or exposing the data to be forgotten. Perceptual quality evaluation further validates our approach’s advantages in both forgetting performance and user experience. These results underscore that our approach not only successfully extends the application boundary of unlearning techniques to source code, but also holds significant implications for advancing reliable code generation.
Index Terms:
Large Language Models, Code Generation, Machine Unlearning
I Introduction
Large language models (LLMs) have revolutionized software engineering, demonstrating remarkable capabilities in various code-related tasks by leveraging training data from massive corpora of open-source code [1, 2, 3]. However, their training data, drawn from broad sources and compiled at massive scale, inevitably contain problematic code, such as copyrighted code files, vulnerable code segments, deprecated APIs, or other undesirable content. With their exceptional memorization capabilities [4], LLMs may retain and reproduce problematic code when assisting users in programming, potentially compromising legal compliance, software security, and code quality. Research [5] has shown that up to 40% of LLM-suggested code may contain security flaws that could lead to serious breaches in production environments. Furthermore, the widespread adoption of LLMs as programming assistants in real-world software development magnifies these concerns, as the issues could lead to significant consequences. Recent high-profile lawsuits against AI companies have drawn considerable attention, with developers challenging AI models for allegedly reproducing proprietary code without permission [6, 7]. This raises a critical question: how can we eliminate the influence of problematic code from LLMs?
Figure 1: Overview of LLM unlearning with the application of removing copyrighted content.
A thorough and direct approach to eliminating the influence of problematic data from LLMs is to retrain LLMs from scratch with clean data. However, this approach incurs prohibitively high computational costs and is impractical for repeated implementation. Recently, LLM unlearning techniques [8, 9, 10, 11, 12] have been proposed to remove the influence of specific training samples while preserving the model’s overall knowledge and functionality, as shown in Figure 1, which holds promise as a potential solution to address the risks posed by problematic data. LLM unlearning techniques leverage post-training strategies to achieve efficient and scalable unlearning without complete model retraining. In natural language processing (NLP) tasks, unlearning techniques have made great progress, such as removing harmful responses, erasing copyrighted content, protecting user privacy, and enforcing policy compliance [9]. However, their application in the domain of code generation remains largely unexplored.
We conduct an empirical study of representative unlearning approaches and their effectiveness in enabling LLMs to forget code. The results show that while these approaches achieve successful forgetting, they often result in significant degradation in model utility. Our analysis reveals that unlearned models often exhibit degraded output, e.g., refusing to generate code, collapsing to high-frequency tokens, producing syntactically incorrect code. We argue that this deterioration stems from the inherent characteristics of code compared to natural language: code exhibits strict syntax, precise dependency modeling, and accurate execution semantics, which are particularly susceptible to disruption. Existing unlearning approaches inadvertently compromise the model’s foundational capabilities in program modeling during the unlearning process. Therefore, It is necessary to develop an unlearning approach applicable to source code to address the problem.
In this paper, we introduce a novel unlearning approach PROD  [^1] PROD stands for Probabilistic Redistribution for Output Distribution of LLMs., which can remove the influence of problematic code from LLMs while effectively preserving the model’s code generation capabilities. PROD optimizes models to fit a target distribution derived from its original output distribution by eliminating the probability of forget data while preserving the distributional structure of the remaining, thereby addressing the model utility degradation of existing unlearning approaches. PROD offers the advantages of lightweight training, efficient optimization, and high usability. We also identify and propose several pressing downstream tasks of LLM unlearning in the domain of source code, including: ❶ copyrighted code unlearning, which enables developers to remove their code contributions from LLMs; ❷ insecure code unlearning, which tackles the issue of LLMs propagating security vulnerabilities; ❸ deprecated API unlearning, which focuses on preventing LLMs from recommending outdated libraries and programming patterns.
Extensive experimental results demonstrate that PROD achieves the best balance between forget quality and model utility among existing LLM unlearning approaches across three downstream tasks. Across four different series of LLMs, our approach consistently demonstrates applicability and effectiveness. To further validate the robustness of PROD, we conduct adversarial attack experiments, and finding that PROD is the only one capable of effectively resisting attacks compared to existing LLM unlearning approaches. Moreover, we conduct perceptual quality evaluation by combining human evaluation and GPT-4 evaluation, and the results shows that PROD has superior forgetting performance coupled with enhanced user experience metrics.
@@ I Introduction
  • We propose a novel LLM unlearning approach PROD, that suppresses the probability of forget data in LLMs’ output distribution while promoting candidate distributional components, effectively addressing the model utility degradation of existing unlearning approaches.

### v2

@@ Abstract
While Large Language Models (LLMs) excel at code generation, their inherent tendency toward verbatim memorization of training data introduces critical risks like copyright infringement, insecure emission, and deprecated API utilization, etc. A straightforward yet promising defense is unlearning, i.e., erasing or down-weighting the offending snippets through post-training. However, we find its application to source code often tends to spill over, damaging the basic knowledge of programming languages learned by the LLM and degrading the overall capability. To ease this challenge, we propose PROD for precise source code unlearning. PROD surgically zeroes out the prediction probability of the prohibited tokens, and renormalizes the remaining distribution so that the generated code stays correct. By excising only the targeted snippets, PROD achieves precise forgetting without much degradation of the LLM’s overall capability. To facilitate in-depth evaluation against PROD, we establish an unlearning benchmark consisting of three downstream tasks (i.e., unlearning of copyrighted code, insecure code, and deprecated APIs), and introduce Pareto Dominance Ratio (PDR) metric, which indicates both the forget quality and the LLM utility. Our comprehensive evaluation demonstrates that PROD achieves superior overall performance between forget quality and model utility compared to existing unlearning approaches across three downstream tasks, while consistently exhibiting improvements when applied to LLMs of varying series. PROD also exhibits superior robustness against adversarial attacks without generating or exposing the data to be forgotten. These results underscore that our approach not only successfully extends the application boundary of unlearning techniques to source code, but also holds significant implications for advancing reliable code generation.
Introduction
Large Language Models (LLMs) have revolutionized the field of software engineering, automatically converting requirements into executable code with efficiency exceeding human programmers (alphacode; Copilot; Agent4code; anthropic_claude_code; jiang2025coderl+). Such a leap is powered by the scaling law, i.e., ever-larger models and ever-growing corpora yield ever-better performance (kaplan2020scaling; hoffmann2022training). However, the training corpora themselves are usually scraped from the open web, almost inevitably containing copyrighted fragments, vulnerable snippets, and deprecated Application Programming Interfaces (APIs), etc. Because LLMs are trained autoregressively, they can memorize and later generate these undesirable snippets (dong-etal-2024-generalization), exposing users to legal risk, security breaches, fragile legacy code, and so on. Empirical studies show that up to 40% of LLM-generated code contains exploitable vulnerabilities (safecoder), while a wave of high-profile lawsuits accuses leading LLMs of violating the copyright of the developers (zhang2024right; butterick2022github).
Purging every problematic code snippet from training corpora and retraining the LLMs from scratch would be straightforward and ideal, yet the cost is obviously unacceptable. A practical alternative is to force the LLMs to selectively forget what it was never meant to know, which is the objective known as unlearning (Unlearning; DBLP:journals/natmi/LiuYJCBHYLXLVBKL25). Given a functional LLM, unlearning algorithms behave like erasers: ❶ they excise the influence of the undesirable samples from the target LLM, while ❷ leaving all other knowledge and capabilities intact. Existing LLM unlearning methods (Unlearning; NPO; DPO; FLAT) typically optimize the model to suppress the likelihood of undesirable outputs by reversing the gradient descent process upon undesirable contents or setting them as negative samples (DBLP:journals/natmi/LiuYJCBHYLXLVBKL25; Unlearning; NPO; DPO; FLAT). e.g., Gradient Ascent (GA) (Unlearning) tunes parameters uphill on samples to be forgotten. For natural language tasks, such methods have been demonstrated effective in excising toxic responses, secret leakage, or policy-violating answers without impairing general capabilities significantly.
Programming languages, however, are bound by rigid syntax, along with strict semantics. When existing unlearning methods directly down-weighting undesirable code snippets, they also strip the language scaffolding that keeps the surrounding code valid. It results in direct utility loss: syntax errors and grammar errors occur, high-frequency tokens repeat, or even the LLMs fall mute. Our pilot empirical study confirms such an issue (see Figure 1): after manipulated by existing unlearning techniques, the target LLMs frequently violate programming language rules or emit nonsensical code snippets. We argue that the root cause here is granularity, i.e., current techniques erase entire slices of knowledge (i.e., the undesirable snippets), damaging the foundational grammar and universal coding patterns the model once mastered through pre-training. Therefore, the open challenge is to eliminate only the offending snippets with surgical precision (maximize forget quality, i.e., the degree that the undesirable contents are eliminated) while leaving model utility (i.e., the general performance of the target LLM) intact.
We introduce Probabilistic Redistribution for Output Distribution (abbreviated as PROD), a surgical, code-oriented unlearning method that forgets code snippets that must be forgotten while leaving all other knowledge of programming languages intact. Unlike existing unlearning approaches, which operate at the granularity of samples, PROD manipulates distributions over individual tokens in a fine-grained granularity. PROD first captures the full distribution over vocabulary predicted by the target LLM, then surgically zeroes out every probability mass assigned to tokens in the undesirable snippet. After pruning incidental noises, it reallocates the probability distribution across the remaining vocabulary in a manner that faithfully preserves statistics of programming languages learned by the target LLM. By optimizing the model to match this carefully sculpted target distribution, PROD attains near-perfect forgetting of the targeted code while effectively maintaining its utility unchanged. We also propose a benchmark for source code unlearning along with a metric, Pareto Dominance Ratio (PDR). The benchmark covers three tasks: ❶ copyrighted code unlearning, where the goal is to remove copyrighted code snippets from LLMs; ❷ insecure code unlearning, which aims to avoid security vulnerabilities produced by LLMs; ❸ deprecated API unlearning, which prevents LLMs from generating APIs from outdated libraries or packages. PDR metric evaluates unlearning approaches by considering both forget quality and model utility.
Our in-depth evaluation over the three unlearning tasks shows that PROD achieves the best PDR score, suggesting its best overall performance between forget quality and model utility. Experiments on four distinct LLMs (i.e., CodeLlama-7B, Qwen2.5-Coder-7B, Deepseek-coder-6.7, Starcoder-7B) demonstrate the broad applicability of PROD. Adversarial attack experiment further reveals that PROD is the only unlearning method among existing LLM unlearning methods resilient against targeted attacks without reproducing forgotten codes.
@@ Introduction
  • We propose a novel unlearning approach PROD, that can precisely forget undesired code snippets and preserve knowledge of programming languages intact in LLMs.